OpenStudy (anonymous):

6.02] For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 9 −x − y = −8 −8 + y −8 − y 8 + y 8 − y

OpenStudy (anonymous):

Does anyone know Algebra?

jimthompson5910 (jim_thompson5910):

What do you get when you solve −x − y = −8 for x ?

OpenStudy (anonymous):

Do you get c?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

it might help to convert −x − y = −8 to x + y = 8 (multiply everything by -1)

OpenStudy (anonymous):

D?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Ok, thx for your help, lol! :D

jimthompson5910 (jim_thompson5910):

hopefully you see how you got it (and it wasn't a guess)?

OpenStudy (anonymous):

Yes, :D I see how to get it, thxxx!

jimthompson5910 (jim_thompson5910):

ok great, yw

OpenStudy (anonymous):

thx agin! Have a good rest of the day!

jimthompson5910 (jim_thompson5910):

you too